10.0 g KCl is dissolved in 1000 grams water. If the density of the solution is 0.997g/ml, calculate:
Molality
Molarity
Mole fraction of KCl
Mass percentage of KCl in the solution
1
Expert's answer
2020-12-10T07:49:11-0500

n=10/74.5 = 0,134228188 moles.

Molality=n/m= 0,134228188/1.01 = 0,132899196 mol/kg.

Molarity = n/V= nxp/m= 0,134228188 x 0.997 / 1.01 = 0,132500498 M.

Mole fraction= 0,134228188/55,6897838 = 0,00241028388.

Mass percentage = 10 x 100%/1010= 0,99009901%.